OpenStudy (anonymous):

Factor 8x^2 + 2x – 3. A. The polynomial is prime. B. (2x – 1)(4x + 3) C. (8x + 1)(x – 3) D. (2x + 3)(4x – 1)

OpenStudy (anonymous):

8*-3=-24 6-4=2 6*-4=-24 2x=(6-4)x=6x-4x write 2x=6x-4x and make factors.

OpenStudy (anonymous):

thx you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

im not getting the right answer